# Internal acoustic resonances: small scattering center compared with the wavelength

### v1

@@ III.2 Systems with a impedance ratio $Z_{1}/Z_{0}\approx 1$
[…]
@@
The results presented in this section underscore the importance of parameter selection in the emergence of resonance effects. The absence of distinct peaks in the energy distribution suggests that moderate impedance differences result in smooth, non-resonant energy behavior in acoustic scattering phenomena.
@@ IV Conclusion
In this study, we present an analytical approach to the single scattering of acoustic waves by spherical objects. We focus on the dynamics inside the scattering center through an energy analysis described by an internal wave, which is influenced by the internal and scattering coefficients. These coefficients exhibit divergences under specific combinations of media and scatterer parameters. These coefficient divergences are associated to wave resonances inside the scatterers. As a result, resonances may occur, even when the scatterer is much smaller than the wavelength of the incident wave, which was not expected. These resonance peaks, driven by the impedance mismatch between the scatterer and the surrounding fluid, lead to substantial energy storage, highlighting the importance of addressing even minor scattering effects to improve system efficiency and reliability. The framework developed here provides a foundation for further investigations into more complex scattering geometries, such as core-shell structures, where interactions between materials could enhance energy localization and control. Another approach is to explore multiple scatterers, non-ideal fluid dynamics, or non-linear effects. These developments would expand the model’s applicability and deepen the understanding of wave scattering across diverse physical contexts. The insights gained from this work contribute to both the theoretical understanding of wave phenomena and practical applications, including the design of acoustic metamaterials [7, 15], soundproofing technologies [16], and virus deactivation [10].

### v2

@@ III.2 Systems with a impedance ratio $Z_{1}/Z_{0}\approx 1$
In this section, we investigate the behavior of both potential and kinetic energies for a system comprising a petrol bubble submerged in seawater. The petrol bubble, with an acoustic impedance of approximately $968\,750\,\text{kg/m}^{2}\cdot\text{s}$, is immersed in seawater, whose impedance is around $1\,558\,528\,\text{kg/m}^{2}\cdot\,\text{s}$. This results in a moderate impedance ratio ($Z_{1}/Z_{0}\approx 0.62$).
Figure 6: Potential energy as a function of $x_{0}$ is given by Eq. (5a), plotted for even ($\ell=0,2,4,6$) and odd ($\ell=1,3,5,7$) modes. The system consists of seawater (medium 0: $\rho_{0}=1024\,\text{kg/m}^{3}$, $c_{0}=1522\,\text{m/s}$) and a petrol bubble (medium 1: $\rho_{1}=968.75\,\text{kg/m}^{3}$, $c_{1}=343\,\text{m/s}$). No significant resonance peaks are observed.
Figure 7: The kinetic energy as a function of $x_{0}$ for the petrol bubble in seawater shows no significant resonance peaks, similar to the potential energy.
@@ IV Conclusion
In this study, we present an analytical approach to the acoustic scattering by spherical objects, based on internal energy analysis. The motivations comes from experiments on isothermal viral inactivation by ultrasound (5–10 MHz), driven primarily by structural resonances in absence of bubble cavitation or related effects. Our findings are consistent with this hypothesis, demonstrating that resonances appear for materials with contrasting acoustic impedance. For materials with similar densities, resonances are supported only if the sound speed of the scattering center is much smaller than that of the surrounding fluid $c_{1}\ll c_{0}$, implying a softer medium with reduced bulk modulus. Advances in acoustic metamaterials have shown that folded materials exhibit reduced sound speeds, owning to their less rigid, spring-like structure, in agreement with the organization of viral nucleoproteins.
Our numerical findings also suggest the emergence of dynamical dissipation mechanisms, even in absence of viscosity, driven by the internal localization of compressional waves. Analogous phenomena are well-documented in optics: once trapped, the waves escape after a characteristic delay $\tau_{\textrm{Wigner}}$, broadening and limiting resonances. As a result, the internal pressure field near resonances receives significant contributions from various modes, with the dominant one characterizing the center of the peak in the small particle regime. This behavior contrasts with the usual monopole-only approximation, which accounts only for radial deformations. Higher-order modes introduce heterogeneous deformations and associated shear forces. In the diffraction regime, even (odd) modes contribute nearly equally, producing a comb-like pattern, confirming that $\ell>0$ must be taken into account. Taken together, monopole resonances in the 1–-10 MHz range are relevant for viral particles of 100–-200 nm, while higher-order modes dominate in the 10–-100 MHz range, indicating two distinct operating regimes. The lower-frequency resonances are fewer and more selective, with their positions strongly dependent on the material properties of the target particle. At higher frequencies, the density of resonances increases, leading to overlapping modes that enhance energy deposition with focus on geometrical aspects.
Our model disregards viscosity as a first approximation, although it is a key aspect in low Reynold flows such as biological systems. Introducing viscosity $\eta$ affects our model in several ways. First, the wave number $k$ becomes a complex to capture the energy loss and a frequency-dependent sound speed $c=c(k)$. Second, the scattering center behaves as a viscoelastic material, capable of both storing and dissipating energy. The internal stress $\sigma_{1}$ acquires a non-trivial spatial distribution, strongly dependent on the velocity field $\mathbf{v}_{1}(\mathbf{r},t)$. In general, the stress decays according to the relaxation time $\tau_{\textrm{relax}}=\eta/E$ where $E$ is the Young modulus. When the stress exceeds the molecular forces that bind them structure together, the medium breaks apart.
The framework developed here provides a foundation for further investigations into more complex scattering geometries, such as core-shell structures, where interactions between materials could enhance energy localization. Another approach is to explore multiple scatterers, non-ideal fluid dynamics, or non-linear effects. These developments would expand the model’s applicability and deepen the understanding of wave scattering across diverse physical contexts. Furthermore, the inclusion of dissipation to our current model would be beneficial for applications in biological settings, where non-inertial forces and stress propagation play crucial roles.